OpenStudy (anonymous):

is the least count of micrometerscrew gauze 0.01mm and vernier calipe 0.1mm?if so,the what is the least count in cms?

OpenStudy (anonymous):

@austinL here's my open q/n!

OpenStudy (austinl):

Just convert those measures that are given in millimeters to centimeters! :D Do you know how to do this?

OpenStudy (anonymous):

no i did it bt m still confused!if u know help me!

OpenStudy (austinl):

Ok, we can use dimensional analysis, are you familiar with this?

OpenStudy (anonymous):

yess a lil'bit!bt plz do help!i will undrstnd!

OpenStudy (austinl):

Ok, here is a site that goes over dimensional analysis. It would behoove you to look through it, as it is a really important concept. Now to apply it to our problem, I will set it up and you can solve, sound good?

OpenStudy (anonymous):

very good!bt where"s the site?

OpenStudy (austinl):

http://www.chem.tamu.edu/class/fyp/mathrev/mr-da.html Silly me, forgot the link. My apologies!

OpenStudy (austinl):

0.01mm Ok, lets convert this to centimeters. We would do so like this! \(\displaystyle \frac{0.01mm}{1}\times\frac{1cm}{100mm}\) You would multiply the top and bottom together and then divide and simplify!

OpenStudy (anonymous):

ok i'll check out!thankyou @austinL :D

OpenStudy (austinl):

Well, top times top, and bottom times bottom to clarify.

OpenStudy (anonymous):

then the answer is 1*10^-4 right?

OpenStudy (austinl):

\(\dfrac{0.01}{100}=0.0001=1\times10^4\) Correctomundo! :D Good job! Do you think you could do it for the second one?

OpenStudy (anonymous):

yess thankyouu for your help!

OpenStudy (austinl):

You are very welcome!!

OpenStudy (anonymous):

nxt tym u need help.tell me !:))

OpenStudy (anonymous):

@austinL @ayumi 1 cm is only 10 mm

OpenStudy (anonymous):

so it's 1 * 10^-3

OpenStudy (austinl):

Oh golly, my bad. I apologize for the error. Good catch @Peter14

OpenStudy (anonymous):

1cm=1mm then how to change 0.01mm to cm?

OpenStudy (anonymous):

@Peter14 help plz!

OpenStudy (anonymous):

1 cm is 10 mm

OpenStudy (anonymous):

then wb 0.01 cm?

OpenStudy (anonymous):

to convert from mm to cm, divide by 10

OpenStudy (anonymous):

ohh thankYou @Peter14